# Review on the decoding algorithms for surface codes

### v4

@@ 5.4.2 TN decoding for the planar code
A surface code encodes a single logical qubit, therefore the normalizer cosists of the following logical cosets: $I_{L}$, $X_{L}$, $Y_{L}$ and $Z_{L}$. As a result of degeneracy, the most probable Pauli error corresponding to a syndrome does not need to correspond to the most probable error class for such syndrome [54]. As mentioned earlier, when facing a non-trivial syndrome in a code, a TN decoder will seek to find a Pauli error correspondent to the syndrome, $E_{rec}$, and the most probable logical coset given the chosen error $E_{rec}$. In this sense, the planar code can be associated to a PEPS tensor network [121] following:
@@
Under these considerations we can define the tensor nodes which form the overall planar code PEPS tensor network in FIG. 18. The elements of the data qubit tensors are the probabilities of experiencing Pauli errors given by the indices. Moreover, the indices $n$, $e$, $s$ and $w$ indicate the tensors to their north, east, south and west; respectively. The indices are binary and indicate if either of the adjacent stabilizer generators operates non-trivially on them as can be seen in the right side of FIG. 18, where one can see that the state of a data qubit tensor is both dependent on the indices and the correspondent Pauli operator on the qubit from the error $E_{rec}$. On the other side, the stabilizer tensor nodes are defined by Kronecker deltas, i.e. they are either 1 when $n=e=s=w$ and 0 otherwise. That is because the stabilizer generators either operate on all their nearest data qubits or do not operate on any of them, as was earlier shown in FIG. 2.
@@
Since the probability of error of data qubits is independent from one to another, considering an error $E_{rec}$ associated to a syndrome and contracting the tensor network from FIG. 17 results in the summation of products of tensor probabilities under all the combinations of the stabilizer generators, which itself is equivalent to finding the probability of the error class $p(\mathcal{G}E_{rec})$. Unfortunately, and as explained before, contracting columns increases the vertical bond dimension exponentially and, thus, the reasonable approach consists in truncating the vertical bond dimension after each contraction to a truncated value $\chi$. This is often done through the Schmidt decomposition [121, 179], which allows the resulting tensor after the column contraction to be represented as a sum of products of smaller tensors called "Schmidt tensors". The SVD is applied to said Schmidt tensors resulting in the Schmidt values, the truncation follows by only considering the "Schmidt tensors" correspondent to the highest $\chi$ Schmidt values [179]. This allows for obtaining approximate values of the probabilities of the error classes, which can then be used for decoding at a higher precision than any of the aforementioned algorithms, if the truncation of the bond dimension is high enough.
@@ 5.4.3 TN decoding for the rotated planar code
Figure 20: Graphical representation of a TN decoding process for a specific syndrome in a 5x5 surface code.
@@ 5.4.4 Performance and threshold
The tensor network decoder has the highest code threshold out of the ones considered in this work equal to $0.185$ under depolarizing noise with a bond dimension $\chi=16$, as can be seen in FIG. 21. It is worth mentioning that this exceptional performance is highly related to the value of $\chi$, the maximum value of which scales in an exponential manner as larger configurations are considered. Recent work has been done studying $\chi$ values for achieving convergence of the tensor network decoding method under several noise model and code tailoring conditions [42, 170].
@@ 5.4.6 Measurement errors
As with the other decoding methods, considering measurement errors significantly increases the complexity of the already complex tensor network decoding method. Nevertheless, the high cost of performing TN decoding allowed for the best performance in the experimental surface code from Google [64].
@@ 5.5.1 Cellular-automaton decoder
Cellular-automaton decoders were first proposed for decoding the toric code [124]. This class of decoders was constructed with the aim of reducing the decoding problem to simple local update rules in the sense that they only depend on the nearest neighbours. The authors justify that a cellular automaton does fit such requirements and, thus, propose the use of an auxiliary cellular automaton to the toric code such that it communicates long-range information between the non-trivial syndromes that have been with the aim that local decision can be made for correcting the associated error. More specifically, the auxiliary cellular automaton stores real numbers $\phi_{t}(x)$ at each time step, $t$, of the decoder for each of its cells. In this sense, the system depends on the error configuration, the activated syndrome elements at a time step and the value of $\phi$, which includes the values for every cell $x$. In order to update the system state from a time $t$ to the next time step $t+1$, $c$ elementary updates of $\phi$ are done based only on the information of the nearest neighbors and another step where the error and the syndrome on the system at such time step is updated too. For doing so, a nearest syndrome is triggered based on an update rule. By repeating this $\tau$ times, the closest potential partners are found when two syndromes cancel each other, implying that a matching has been found, recovery Pauli operation. The authors in [124] propose that phiphi represents an attractive force between the measured syndrome, emulating an electrostatic or gravitational field which can be discretized into update rules since they can be discretized to local dynamics. Therefore, they propose to emulate the long field attraction of those fields to represent that the closest triggered syndromes are attracted, and they propose the local dynamics induced by those laws to updated the values locally.
Figure 22: Visual representation of the 2D cellular automaton decoder for the toric code with periodic boundaries. The blue dots, represented by lattice $V$, are the data qubits while the green boxes, represented by lattice $\Lambda$, depict the cellular automaton decoder. Figure taken from [124].
@@
The update rule proposed by the authors is based on selecting which of the neighbour syndromes is activated as a function of the field intensity provided by the values of $\phi$. In this sense, they look for the the adjacent cell with largest field value and then move there with a probability $1/2$. Regarding the function $\phi$, a local discretization of the Gauss’ law was selected for emulating the long field attraction [124]. There have been other proposals for the update rule of the cellular automaton decoder such as the so called Toom’s rule [183, 184] or the sweep rule [185, 186]. Most of the cellular automaton decoders have been studied for toric codes with boundaries in $d$ dimensions [124, 183, 184, 185, 186], but some lattices without boundaries have examined [186]. Additionally, it was shown that the decoders for the toric code can be applied to decode color codes [187] and, thus, cellular automaton decoders can also be applied for this family of QECCs. Unfortunately, its application for a rotated planar code has not been investigated.
@@
Regarding the threshold achieved by this class of decoders it is noteworthy to say that the noise model considered has been primarily an i.i.d. phase flip model [124, 183, 184, 185, 186]. In this sense, a threshold of $8.2\%$ was claimed in [124] for a 2D toric code, while a MWPM decoder exhibits a $10.31\%$ threshold for such code [188]. However, a threshold of $15.5\%$ was claimed in [184] for the cubic toric code using the sweep rule. In terms of the complexity of this decoder, it can naturally be integrated in a parallel way by communicating the processing cores locally and, therefore, it is usually seemed as a fast decoder [124, 185, 186]. In this sense, the 2D cellular automaton of [124] requires a number of updates in the order of $\log^{5}(d)$ updates, while the 3D implementation requires a number of updates in the order of $\log^{3}(d)$. As the authors discuss, this comes from the fact that the cellular automata decoder accepts a parallel implementation from construction [124]. The sweep decoder implementation in [185] requires a runtime $\mathcal{O}(d)$, so linear in $d$.
@@
An important feature of cellular automaton decoders is that they are robust for the case that measurement errors are considered [185, 186]. Specifically, this decoder is deemed to be a single-shot decoder, which implies that measurement errors can be dealt with a single decoding round (recall Appendix B). This is specially interesting for a decoder since as discussed before, the way in which decoders usually deal with noisy syndromes is by performing $d$ measurement rounds and then solving the space-time like obtained graph [186].
@@ 5.5.2 Renormalization group decoder
The renormalization group (RG) decoder was proposed as a fast and efficient decoder for the Kitaev toric code [123]. The main idea behind this decoder is that the toric code can be seen as a concatenated code so that the decoding problem can be resolved from the smaller codes that form it. The toric code is not exactly a concatenated code, but the authors propose that it can be actually considered to be one when allowing the blocks of the concatenation to share qubits, so they are actually overlapped [123]. Following this rationale, the RG decoder computes the logical error probabilities of the smallest codes in the concatenation, which then are used as the noise model for the second layer of codes on the concatenation. the process is repeated until reaching the top layer of the code, which actually reperesents the complete toric code and, thus, the estimation of the error is obtained for correction. The authors discuss that since the small codes for the concatenation are small open-boundary topological codes, those can be decoded by brute force [123].
@@
As stated, the renormalization decoder was originally proposed for the Kitaev toric code, so with periodic boundaries, but modifications for decoding color codes [189], the 3D cubic code [190, 191, 192] qudit topological codes [191, 193, 194] and the four-dimensional toric code (both for periodic and open boundary conditions) [195] have also been proposed. Interestingly, the RG decoder was the first known algorithm to decode the color code. The RG decoder has not been investigated for the rotated planar code discussed thorugh this paper, but it could be valid, in principle, by using the ideas for decoding open boundary four-dimensional toric codes [195] since the rotated planar code is essentially a rotated toric code with boundaries.
Figure 23: Example of a code block decomposition used for RG decoding. Each of the edges represnt a qubit (12 total). In this example, a subcode of two qubits is used for RG. The dashed lines represent qubits shared by two blocks due to the fact that the toric code is not exactly a concatenated code. Figure taken from [123].
@@
Regarding the code threshold achieved by RG decoders, its standard version shown a threshold of $7.8\%$ for the depolarizing noise model, which falls short when compared to the $15.5\%$ threshold obtained by decoding it with the MWPM decoder. However, the application of belief-propagation decoder before the RG decoder was proposed in [123]. This was done so that the smallest codes of the concatenation can start the decoding with the probability distribution given by BP implying that the RG is more accurate globally. By doing so, the BP+RG decoder achieved a $16.5\%$ threshold, which exceeds that of the MWPM decoder by itself [123]. This, however, was comparing the performance with the MPWM alone, i.e. belief-matching would also increase the threshold [158, 87]. In addition, it does not exceed the threshold achieved by the tensor network decoder, which stands between $17\%$ and $18.5\%$ [121].
@@
Similar to the cellular automata decoder, the RG decoder admits a parallelization of the decoding problem since each of the codes forming a layer of the concatenation can be decoded in at the same time [123]. This implies that the complexity of the algorithm is a function of the number of concatenation layers, which scales logarithmically with the distance and, thus, number of qubits of the code. Therefore, the decoding complexity of the standard RG decoder is in the order of $\mathcal{O}(d)$ if the complete parallelization is done, while in the order of $\mathcal{O}(d^{2}\log{d})$ if the decoding is done serially [123]. Therefore, this decoder is considered to be a fast decoder. The BP+RG approach for increasing the threshold will suffer a toll in the complexity coming from having to execute BP too. It is noteworthy to say that whenever faulty syndrome measurements coming from circuit level noise are considered, the RG decoder needs $d$ rounds of measurements for dealing with them, implying that the complexity of the decoder will also increase [186].

### v5

@@ 5.4.2 TN decoding for the planar code
As mentioned earlier, when facing a non-trivial syndrome in a code, a TN decoder will seek to find a Pauli error correspondent to the syndrome, $E_{rec}$, and the most probable logical coset given the chosen error $E_{rec}$. In this sense, the planar code can be associated to a PEPS tensor network [60] following:
@@
Under these considerations we can define the tensor nodes which form the overall planar code PEPS tensor network in FIG. 19, which will be defined by the following expressions corresponding to the tensor figures from top to bottom:
$H^{i}_{n,e,s,w}$ $=p(E^{i}_{rec}+(nZ)+(eX)+(sZ)+(wX)),$ (34)
$V^{i}_{n,e,s,w}$ $=p(E^{i}_{rec}+(nX)+(eZ)+(sX)+(wZ)),$ (35)
$s^{i}_{n,e,s,w}$ $=\delta_{n,e,s,w}.$ (36)
The elements of the data qubit tensors are the probabilities of experiencing Pauli errors given by the indices. Moreover, the indices $n$, $e$, $s$ and $w$ indicate the tensors to their north, east, south and west; respectively. The indices are binary and indicate if either of the adjacent stabilizer generators operates non-trivially on them as can be seen in the right side of FIG. 19, where one can see that the state of a data qubit tensor is both dependent on the indices and the correspondent Pauli operator on the qubit from the error $E_{rec}$. On the other side, the stabilizer tensor nodes are defined by Kronecker deltas, i.e. they are either 1 when $n=e=s=w$ or 0 otherwise. That is because the stabilizer generators either operate on all their nearest data qubits or do not operate on any of them, as was earlier shown in FIG. 2.
@@
Since the probability of error of data qubits is independent from one to another, considering an error $E_{rec}$ associated to a syndrome and contracting the tensor network from FIG. 18 results in the summation of products of tensor probabilities under all the combinations of the stabilizer generators, which itself is equivalent to finding the probability of the error class $p(\mathcal{L}E_{rec})$. Unfortunately, and as explained before, contracting columns increases the vertical bond dimension exponentially and, thus, the reasonable approach consists in truncating the vertical bond dimension after each contraction to a truncated value $\chi$. This is often done through the Schmidt decomposition [60], which allows the resulting tensor after the column contraction to be represented as a sum of products of smaller tensors called "Schmidt tensors". The SVD is applied to said Schmidt tensors resulting in the Schmidt values, the truncation follows by only considering the "Schmidt tensors" correspondent to the highest $\chi$ Schmidt values. This allows for obtaining approximate values of the probabilities of the error classes, which can then be used for decoding at a higher precision than any of the aforementioned algorithms, if the truncation of the bond dimension is high enough.
@@ 5.4.3 TN decoding for the rotated planar code
Figure 21: Graphical representation of a TN decoding process for a specific syndrome in a distance-5 rotated code.
@@ 5.4.4 Performance and threshold
The tensor network decoder has the highest code threshold out of the ones considered in this work equal to $0.185$ under depolarizing noise with a bond dimension $\chi=16$, as can be seen in FIG. 22. It is worth mentioning that this exceptional performance is highly related to the value of $\chi$. Recall that $\chi$ is the truncation is the maximum number of Schmidt tensors that we will be considered when contracting the code. A larger distance will increase the number of Schmidt tensor exponentially, thus, the maximum value of $\chi$ scales in an exponential manner as larger code distances are considered. Recent work has been done studying $\chi$ values for achieving convergence of the tensor network decoding method under several noise model and code tailoring conditions [33, 141].
@@ 5.4.6 Measurement errors
As with the other decoding methods, considering measurement errors significantly increases the complexity of the already complex tensor network decoding method. Nevertheless, the high cost of performing TN decoding allowed for the best performance in the experimental rotated code from Google [51].
@@ 5.5.1 Cellular-automaton decoder
Before the cellular-automaton decoder is introduced, we introduce the concept of cellular-automaton. A cellular automaton [146] can be described as a set of cells in a grid, each of which is described by one of a finite number of states. For each cell, a set of cells called the neighbourhood is defined relative to it. An initial state for all the cells is selected at time instance $t=0$. Then, a new configuration is generated according to a fixed rule dependent to the state of each cell and the state of the cells in its neighbourhood, advancing $t$ by 1. Thus, the overall system evolves in time following its initial state and the fixed rule.
An example of a cellular automaton is the “Game of Life" [147]. In the Game of Life there is a square grid in which the neighbourhood of every cell is composed by its left, right, top, bottom and diagonal cells. For the game of life, each cell can take two values, "alive" or "dead". Moreover, given a time $t+1$, the updated state of every cell will be given by the following rules:
  • Any alive cell with fewer than two alive neighbours dies.
  • Any alive cell with two or three alive neighbours continues alive.
  • Any alive cell with more than three alive neighbors dies.
  • Any dead cell with exactly three alive neighbours becomes an alive cell.
An instance of this cellular-automaton is illustrated in FIG. 24. The right alive cell in the top figure dies because it is only adjacent to a single alive cell and the top-left cell becomes alive in the second plot because it is adjacent to three alive cells. The remaining stay alive because they have 2 or 3 alive adjacent cells.
Figure 23: Illustration of two time instances in the Game of Life. On the top figure, an illustration of a configuration in time $t$, where violet squares are live cells and white ones are dead. On the bottom figure, the configuration of the cellular automaton at time instance $t+1$ is presented.
Cellular-automaton decoders were first proposed for decoding the toric code [100]. This class of decoders was constructed with the aim of reducing the decoding problem to simple local update rules. The authors justify that cellular automata do fit such requirements and, thus, propose the use them in order to decode toric codes.
The approach on the first cellular-automaton decoder consists in considering two cellular automata, a main one and an auxiliary one, where cells the states of the checks given an error [100]. Within the main cellular automaton each cell can be in one of two states, trivial or non-trivial. FIG. 24 provides a portrayal of the mapping of a toric code under a syndrome to its cellular automaton representation.
Figure 24: Illustration of the mapping of a syndrome in the toric code to a the check cellular automaton. In the left, a distance-4 toric code with a syndrome, where non-trivial syndrome elements are denoted with exclamation marks. To the right, the cellular-automaton to consider, where green cells indicate trivial checks and red ones indicate non-trivial checks.
The cellular-automaton decoder considers that every non-trivial cell $x$ corresponds to a charge $q(x)$. Moreover, an electrostatic field $\phi(x)$ is considered within every configuration of the cellular automaton. The fixed rule that characterizes the main cellular automaton system is that on every time instance, non-trivial cells consider their neighbourhood, and move with a probability of $1/2$ to the adjacent cell with highest field value, otherwise, they remain in the same spot as in the previous instance. Whenever two non-trivial cells collide they disintegrate and the decoder considers them as matched. The field is then changed following the next cellular automaton configuration. The cellular-automaton decoder will terminate when all non-trivial checks have been disintegrated and all cells are trivial. At the end, the recovered error will correspond to operators matching collided pairs.
The field utilized by the main cellular automaton in each iteration is provided by the auxiliary cellular automaton, which, for this general explanation, will take the same lattice form as the main cellular automaton. In the first time instance $t=0$, every cell within the auxiliary cellular automaton will have a value $q_{E}(x)$, which will be $q$ or 0 if they correspond to a non-trivial or trivial cell respectively. Afterwards, for every time instance, the state of the auxiliary cellular automaton cells will evolve as:
$\phi_{t+1}(x)=(1-\eta)\phi_{t}(x)+\frac{\eta}{2D}\sum_{\langle y,x\rangle}\phi_{t}(y)+q_{E}(x),$ (37)
where $\phi_{t}(x)$ indicates the field in the cell $x$ and the time instance $t$, $0<\eta<1/2$ is a smoothing parameter we can freely choose, $D$ is the dimension of the auxiliary cellular automaton lattice, which, for our case, would be 2 and the sum $\langle y,x\rangle$ is over all the cells in the neighbourhood of the $x$ cell. In order to set the field for the main cellular automaton, $c$ evolution instances are contemplated on the auxiliary cellular automaton. The $c$ parameter is tunable and known as the field velocity and needs to grow with logarithmic dependence with the size of the auxiliary matrix in order to establish meaningful fields for the main cellular automaton [100].
@@
There have been other proposals for the update rule of the cellular automaton decoder such as the so called Toom’s rule [148, 149] or the sweep rule [150, 151] altogether with other lattice structures for the auxiliary cellular automaton such as 3D toric codes [100]. Most of the cellular automaton decoders have been studied for toric codes with boundaries in $d$ dimensions [100, 148, 149, 150, 151], but some lattices without boundaries have examined [151]. Additionally, it was shown that the decoders for the toric code can be applied to decode color codes [152] and, thus, cellular automaton decoders can also be applied for this family of QECCs. Unfortunately, its application for a rotated planar code has not been investigated.
@@
Regarding the threshold achieved by this class of decoders it is noteworthy to say that the noise model considered has been primarily an i.i.d. phase-flip model [100, 148, 149, 150, 151]. In this sense, a threshold of $8.2\%$ was claimed in [100] for a 2D toric code, while a MWPM decoder exhibits a $10.31\%$ threshold for such code [153]. However, a threshold of $15.5\%$ was claimed in [149] for the cubic toric code using the sweep rule. In terms of the complexity of this decoder, it can naturally be integrated in a parallel way by communicating the processing cores locally and, therefore, it is usually seemed as a fast decoder [100, 150, 151]. In this sense, the 2D cellular automaton of [100] requires a number of updates in the order of $\log^{5}(d)$ updates, while the 3D implementation requires a number of updates in the order of $\log^{3}(d)$. As the authors discuss, this comes from the fact that the cellular-automaton decoder accepts a parallel implementation from construction [100]. The sweep decoder implementation in [150] requires a runtime $\mathcal{O}(d)$, so linear in $d$.
@@
An important feature of cellular-automaton decoders is that they are robust for the case that measurement errors are considered [150, 151]. Specifically, this decoder is deemed to be a single-shot decoder, which implies that measurement errors can be dealt with a single decoding round (recall Appendix B). This is specially interesting for a decoder since as discussed before, the way in which decoders usually deal with noisy syndromes is by performing $d$ measurement rounds and then solving the space-time like obtained graph [151].
@@ 5.5.2 Renormalization group decoder
[…]
Figure 25: Graphical depiction of $n$ layers of code concatenation. For this example, the same code is repeated every layer. Such code encodes a single qubit within three qubits. Figure taken from a talk by Duclos-Cianci (Accessible in: https://av.tib.eu/media/35296).
FIG. 25 shows a graphical representation of code concatenation. Decoding codes that present this concatenated structure can be done by decoding each of the layers, which consists of several smaller codes. Specifically, the process begins by decoding each of the small codes of the last layer of concatenation (considering the syndrome information of such layer) and other than taking a hard decision for the logical error, the soft information of the logical error probability is passed to the next layer as a priori information. This process is repeated until the last layer is reached, where the hard decision for the estimated logical error is done. For the first decoding step, the a priori information is the physical error probability. FIG. 26 shows this process graphically. Importantly, recall that the optimal decoding problem, i.e. considering degeneracy, for stabilizer codes is not efficiently solvable when the code size grows. However, it can actually be dealt with efficiently for small codes by brute force. Therefore, codes with a concatenated structure as the one presented in FIG. 25 can actually be optimally decoded by following the described process.
Figure 26: Graphical depiction of the decoding process for the concatenated code in FIG. 25. The process goes from the bottom to the top. The red lines represent the fact that the logical error probability is passed to the next layer as apriori information for decoding. The $L$ operator in the top layer represents the estimated logical error. Figure taken from a talk by Duclos-Cianci (Accessible in: https://av.tib.eu/media/35296).
The proposal of the RG decoder for the toric code was motivated by such ideas [99]. The toric code is not exactly a concatenated code, but the authors propose that it can be approximated as one when allowing the blocks of the concatenation to share qubits, so the the blocks are actually overlapped [99]. Each of the small codes in this approximated concatenation is selected to be an small (12 qubit) open boundary toric code. This unit cell is shown in FIG. 27. In the unit cell depicted there, the dashed data qubits represent shared qubits between neighbouring unit cells in the periodic structure of the toric code. In addition, the measured syndromes of the filled checks are used in the decoding of the unit cells, while the empty check information is used to decode the next concatenation step (as seen in the right figure filled checks). Following this rationale, the RG decoder computes the logical error probabilities of the smallest codes in the concatenation, which then are used as the noise model for the second layer of codes on the concatenation. Since the concatenated structure is an approximation, the soft information obtained for the shared qubits is different in each of the blocks. This implies that the marginal probabilities of error obtained after decoding differ, i.e.
$P_{L,i}(E)\neq P_{R,i}(E),$ (40)
where $i$ refers to one of the shared qubits, $L$ and $R$ refers to let and right cells and $E$ the error experienced. To solve this inconsistency, the authors proposed to use belief propagation to obtain the merged soft information from the independently obtained ones and, thus, force the marginal to be equal, i.e. $P_{L,i}(E)=P_{R,i}(E)$. FIG. 28 shows a graphical depiction of two neighbouring unit cells for which shared qubits consistency is required, and obtained via BP. This process is done, as described for concatenated coding before, until the last layer of concatenation is reached and the estimation of the error is obtained for correction. The authors discuss that since the small codes for the concatenation are small open-boundary topological codes, those can be decoded by brute force [99].
Figure 27: Example of a code block decomposition used for RG decoding. The unit cells consist of 12 data qubits. The data qubit and check color convention is the same as through the rest of the article. Dashed data qubits refer to qubits shared between unit cells. Filled checks are used in decoding each layer while empty ones are passed to the next concatenation layer. Since this code block is an open boundary toric code, two logical qubits are encoded as depicted in the right part of the figure. Note also that here the information of the unused checks of the left is used. Figure inspired from [99].
@@
As stated, the renormalization decoder was originally proposed for the Kitaev toric code, so with periodic boundaries, but modifications for decoding color codes [154], the 3D cubic code [155, 156, 157] qudit topological codes [156, 158, 159] and the four-dimensional toric code (both for periodic and open boundary conditions) [160] have also been proposed. Interestingly, the RG decoder was the first known algorithm to decode the color code. The RG decoder has not been investigated for the rotated planar code discussed through this paper, but it could be valid, in principle, by using the ideas for decoding open boundary four-dimensional toric codes [160] since the rotated planar code is essentially a rotated toric code with boundaries.
Figure 28: Example of two unit cells of the toric code decomposition where shared qubits exchange information by means of belief propagation to reach consistency.
@@
Regarding the code threshold achieved by RG decoders, its standard version shown a threshold of $7.8\%$ for the depolarizing noise model, which falls short when compared to the $15.5\%$ threshold obtained by decoding it with the MWPM decoder. However, the application of belief propagation decoder before the RG decoder was proposed in [99]. This was done so that the smallest codes of the concatenation can start the decoding with the probability distribution given by BP implying that the RG is more accurate globally. By doing so, the BP+RG decoder achieved a $16.5\%$ threshold, which exceeds that of the MWPM decoder by itself [99]. This, however, was comparing the performance with the MPWM alone, i.e. belief-matching would also increase the threshold [65, 133]. In addition, it does not exceed the threshold achieved by the tensor network decoder, which stands between $17\%$ and $18.5\%$ [60].
@@
Similar to the cellular automata decoder, the RG decoder admits a parallelization of the decoding problem since each of the codes forming a layer of the concatenation can be decoded in at the same time [99]. This implies that the complexity of the algorithm is a function of the number of concatenation layers, which scales logarithmically with the distance and, thus, number of qubits of the code. Therefore, the decoding complexity of the standard RG decoder is in the order of $\mathcal{O}(d)$ if the complete parallelization is done, while in the order of $\mathcal{O}(d^{2}\log{d})$ if the decoding is done serially [99]. Subsequently, this decoder is considered to be a fast decoder. The BP+RG approach for increasing the threshold will suffer a toll in the complexity coming from also having to execute BP. It is noteworthy to say that whenever faulty syndrome measurements coming from circuit level noise are considered, the RG decoder needs $d$ rounds of measurements for dealing with them, implying that the complexity of the decoder will also increase [151].
Figure 29: Example of decomposing a Pauli error (left), $E$, into its stabilizer (middle, left), $S$; pure error (middle,right), $T$; and logical error (left), $L$; components as in equation (41). The green squares are depicted to show that such Pauli chain forms an stabilizer for the $S$ component. Note that the pure error is the only Pauli string triggering syndromes. Figure motivation taken from [101].